OpenStudy (anonymous):

How could I find x: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (anonymous):

\[-4(x+10)-6=-3(x-2)\]\[-4x-40-6=-3x+6\]\[-4x-46=-3x+6\]\[-x=52\]\[x=-52\]

OpenStudy (anonymous):

Thanks

OpenStudy (noseboy908):

Also, be sure to close the question if it's been answered. It helps keep the open questions column less cluttered.